# Binarized Simplicial Convolutional Neural Networks

### v1

@@ Abstract
Graph Neural Networks have a limitation of solely processing features on graph nodes, neglecting data on high-dimensional structures such as edges and triangles. Simplicial Convolutional Neural Networks (SCNN) represent higher-order structures using simplicial complexes to break this limitation albeit still lacking time efficiency. In this paper, we propose a novel neural network architecture on simplicial complexes named Binarized Simplicial Convolutional Neural Networks (Bi-SCNN) based on the combination of simplicial convolution with a binary-sign forward propagation strategy. The usage of the Hodge Laplacian on a binary-sign forward propagation enables Bi-SCNN to efficiently and effectively represent simplicial features that have higher-order structures than traditional graph node representations. Compared to the previous Simplicial Convolutional Neural Networks, the reduced model complexity of Bi-SCNN shortens the execution time without sacrificing the prediction performance and is less prone to the over-smoothing effect. Experimenting with real-world citation and ocean-drifter data confirmed that our proposed Bi-SCNN is efficient and accurate.
@@
Graph Learning; Graph Neural Networks; Simplicial Complex; Convolutional Neural Networks
@@ 1 Introduction
Graphs and networks are gaining enormous awareness in the field of machine learning and artificial intelligence owing to their ability to effectively learn and represent irregular data interactions [1, 2, 3]. Graph structures can be deduced from various real-life applications by embedding the features onto the nodes of graph representations using the graph Laplacian. Sensor graphs of nationwide temperature [4] and air quality [5] are built by combining sensor readings with the sensor location; citation graphs [6] can be formed by linking the citation relationship between authors. In the past, spatial and spectral graph neural networks (GNNs) such as the Graph Convolutional Networks (GCN) [7, 8], Graph Attention Networks (GAT) [9], Graph SAGE [10] and their variants have had major success on graph representation learning tasks including node classification, graph classification, and link prediction. The GCN and the GAT can be generalized by the Message Passing Neural Networks where instead of Global level graph representation done by the graph Laplacian matrix or the graph Adjacency matrix, a localized message-passing scheme is defined for each node [11]. However, GNNs operate only on the features residing on the graph nodes and have not given enough attention to the graph edges and the structures formed by the edges. In the real world, the fact is that data or features can appear not only on the graph nodes but also on higher-order structures. Traffic flow could be recorded on either the edges [12] or the nodes [13] of a graph. A map of ocean drifter trajectories can be modeled using the data on graph edges [14]. In a citation complex, a collaboration among three or even more authors is modeled using simplicial complexes, which is a scenario that cannot be simply recorded using the graph nodes only because a traditional representation of an edge connection between two nodes can only represent a one-to-one relationship between two authors [15]. More such examples where data appears on structures with higher-order than graph nodes include pandemic prediction based on population mobility among regions [16] and dynamic gene interaction graphs [17].
@@
Recently the concept of Topological Signal Processing was introduced: instead of the graph Laplacian, a more generalized Laplacian known as the Hodge Laplacian is adopted in Topological Signal Processing to incorporate the representation of higher-order features that are beyond the graph nodes [18]. In Topological Signal Processing, operations such as spectral filtering, spatial diffusion, and convolution are defined on simplicial complexes using the Hodge Laplacians [19]. High-dimensional feature representations such as data on the graph edges and data on the triangles formed by three edges can be embedded into simplicial complexes using methodologies analogous to graph (node) embedding [20].
@@
If we feed the outputs of the Topological Signal Processing into non-linear active functions the concatenate a few of these together, we can transform them into neural networks. The Simplicial Neural Network (SNN) generalized the GCN onto simplicial complexes by using Topological Signal Processing to extend the convolutional ability of GCN to process higher-order simplicial features [15]. The Simplicial Convolutional Neural Networks (SCNN) [21] further improves the performance of SNN by applying the Hodge decomposition to the simplicial convolution so it is able to train the upper and lower simplicial adjacencies separately. The Simplicial Attention Network (SAT) [22] and the Simplicial Attention Neural Network [23] introduced the attention mechanism to SNN and SCNN to process simplicial features, making them generalized versions of the GAT from graphs to simplicial complexes. The Simplicial Message Passing Networks have generalized the message passing on graphs [11] to simplicial complexes [24]. The Dist2Cycle was proposed to use Simplicial complexes to define neural networks for Homology Localization [25]. These architectures have shown to be successful at applications such as imputing missing data on the citation complexes of different data dimensions and classifying different types of ocean drifter trajectories.
@@
Although the simplicial complex-based neural networks have achieved numerous successes, the focus of the previous networks was on how to extend concepts such as convolution or attention from graphs onto simplicial complexes, which leads to two drawbacks. First, the recently proposed simplicial neural network structures do not prioritize time efficiency, resulting in inefficiency in terms of computational complexity and run time. Let us look into the GCN, which is the 0-simplex analogy of the SNN, the Simplified GCN [26] has proved a trained GCN essentially is a low-pass filter in the spectral domain and the complexity of GCN can be reduced by eliminating nonlinearities and collapsing weights. The Binary-GCN (Bi-GCN) [27] was inspired by the Binary Neural Networks to use binary operations such as XOR and bit count to speed up training time. In the field of Graph Signal Processing, the Adaptive Graph-Sign algorithm derived from $l_{1}$-optimization results in a simple yet effective adaptive sign-error update that has outstanding run speed and robustness [28, 29]. To the best of our knowledge, similar improvements that focus on the efficiency and complexity of neural network algorithms on the simplicial complex are absent. Following the first drawback of the current neural network algorithms on simplicial complexes, the high complexity of current simplicial neural network structures may cause the resulting simplicial embeddings to get over-smoothed in some scenarios. In their graph predecessors, the over-smoothed phenomenon was extensively analyzed [30, 31] but has not yet gained enough attention in the simplicial complex-based neural networks.
@@
    The Topological Signal Processing foundation of the Bi-SCNN enables Bi-SCNN to process data that have higher-order structures than graph nodes, allowing the Bi-SCNN to effectively represent simplicial data features.
@@
    The Bi-SCNN can be interpreted using spatial and spectral filters defined using the Topological Signal Processing backbone.
@@
    The combination of simplicial convolution with a binary-sign forward propagation strategy significantly reduces the model complexity, leading to a reduced training time when compared to existing algorithms such as SCNN, without sacrificing the prediction accuracy. Additionally, the formulation of the Sign() function in the binary-sign forward propagation makes it naturally an activation function.
@@
    Bi-SCNN is less prone to over-smoothing as a result of the reduced architectural complexity.
@@ 2 Related Works
Algorithms that conduct convolution on the nodes of a graph such as GCN [8] operate only on the graph nodes, but not on $k$-simplexes with $k\geq 1$. Both the SNN [15] and the SCNN [21] are direct extensions of the GCN onto simplicial complexes; SNN is architecturally identical to the SCNN except SNN does not use Hodge decomposition. The Bi-SCNN is an extension of the SCNN [21], but the Bi-SCNN uses feature normalization and feature binarization which the SCNN does not. Also, the binary-sign forward propagation strategy in the Bi-SCNN layer has 2 outputs but the SCNN layer only has 1 output, making a neural network that contains more than 2 layers of Bi-SCNN a drastically different architecture than using purely SCNN.
@@
The Bi-GCN [27] has an acronym similar to the Bi-SCNN, but we should emphasize that the Bi-SCNN is not simply a simplicial complex version of the Bi-GCN and the two are drastically different from each other with the following four differences. Firstly, the Bi-GCN operates only on the graph nodes, while the Bi-SCNN operates on $k$-simplexes. Secondly, as explained earlier, the Bi-SCNN does not follow the traditional GCN/SCNN forward propagation strategy because the feature binarization and the feature normalization propagate separately. Thirdly, the Bi-GCN binarizes both the feature and the trainable weights but Bi-SCNN only binarizes the features. Fourthly, the Bi-GCN uses binary operations bit count and XOR to speed up the same GCN forward propagation and uses predefined values to deal with the backpropagation of binary operations; the Bi-SCNN uses a binary-sign forward propagation and we do not make any predefinitions for backpropagation.
3 Background
@@
We will begin with a set of vertex $\mathcal{V}={\{v_{0}...v_{N_{0}}}\}$ to define a $k$-simplex $\mathcal{S}_{k}$, where $\mathcal{S}_{k}$ is a subset of $\mathcal{V}$ with cardinality $k+1$. Throughout the paper, subscript $k$ is used to denote which $\mathcal{S}_{k}$ a variable belongs to. The collection of $k$-simplexes with $k=0...K$ is a simplicial complex $\mathcal{K}$ of order $K$. If generalizing a graph $\mathcal{G}=\{\mathcal{V,E}\}$ as a simplicial complex, the vertex set $\mathcal{V}$ of $\mathcal{G}$ is the node set that belongs to $\mathcal{S}_{0}$ and the edge set =$\mathcal{E}=\{e_{1}...e_{N_{1}}\}$ belongs to $\mathcal{S}_{1}$. Note that for a simplex $\mathcal{S}_{k}\in\mathcal{K}$, if $\mathcal{S}_{k-1}$ is a subset of $\mathcal{S}_{k}$, then $\mathcal{S}_{k-1}\in\mathcal{K}$. We use the variable ${N_{k}}$ to denote the number of elements of $\mathcal{S}_{k}$. Take a graph $\mathcal{G}$ as an example, ${N_{0}}=|\mathcal{V}|$ is the number of nodes and ${N_{1}}=|\mathcal{E}|$ is the number of edges. Each simplicial feature vector of a simplex $\mathcal{S}_{k}$ is a mapping from the $k$-simplex to $\mathbb{R}^{N_{k}}$ and will be denoted as $\boldsymbol{x}_{k}=[x_{k,0}...x_{k,N_{k}}]^{T}$ [18] . Simplicial data with $d$ features will be denoted as $\mathbf{X}_{k}$ = [$\boldsymbol{x}_{k,0}...\boldsymbol{x}_{k,d}$].
@@
One can see an important characteristic of $\mathbf{L}_{k}$, for $1\leq k<K$, $\mathbf{L}_{k}$ can be split into $\mathbf{L}_{k,l}$ and $\mathbf{L}_{k,u}$ (1); $\mathbf{L}_{k,l}$ is the lower Hodge Laplacian that represents the lower adjacencies of $\mathcal{S}_{k}$, and $\mathbf{L}_{k,u}$ is the upper Hodge Laplacian that represents the upper-adjacency of $\mathcal{S}_{k}$ [18]. It is worth pointing out an important property of the SFT in the spectral domain: the eigenvectors paired with the non-zero eigenvalues of $\mathbf{L}_{k,u}$ are orthogonal to the eigenvectors paired with the non-zero eigenvalues of $\mathbf{L}_{k,l}$ [18].
The graph Laplacian matrix seen in GCN or GAT is $\mathbf{L}_{0}$. Two examples of incidence matrices are $\mathbf{B}_{1}$ and $\mathbf{B}_{2}$, which are the node-to-edge incidence matrix and the edge-to-triangle incidence matrix respectively.
4 Derivations of the Binarized Simplicial Convolutional Neural Networks
@@
We will begin with the simplicial convolution on a $k$-simplex, which is the theoretical foundation of our Bi-SCNN, along with a discussion of spectral and spatial Topological Signal Processing techniques. Then, we will provide the derivations of Bi-SCNN and combine binary-sign updates, simplicial filters, and nonlinearities to form our Bi-SCNN layer. The goal is to improve the simplicial convolution shown in (7) by reducing its computational complexity and making it less prone to over-smoothing.
4.1 Simplicial Convolution
In GNNs, operations such as graph convolution and embedding of the node features $\mathbf{X}_{0}$ are possible by the computations defined by the graph Laplacian. For instance, the spatial graph convolution in GCN aggregates the data from the 1-hop neighbors of each node by $\mathbf{L}_{0}\mathbf{X}_{0}$ [8]. Graph convolution can also be defined by applying low-pass filters in the spectral domain to data on the graph nodes [7].
In Topological Signal Processing, an equivalent analogy can be drawn by defining spatial and spectral operations using the Hodge Laplacian. The convolution is essentially transforming spatial data into the spectral domain, applying a filter in the spectral domain to the simplicial data, and using the inverse transform to transform the filtered data back into the spatial domain [18]. For a set of features to be able to transform between domains in a $k$-simplex, we need the simplicial Fourier transform (SFT) defined by the eigendecomposition of the Hodge Laplacian matrix. For the $k$-simplex, the SFT is the eigendecomposition

### v2

@@ Abstract
Graph Neural Networks have the limitation of processing features solely on graph nodes, neglecting data on high-dimensional structures such as edges and triangles. Simplicial Convolutional Neural Networks (SCNN) represent high-order structures using simplicial complexes to break this limitation but still lack time efficiency. In this paper, a novel neural network architecture named Binarized Simplicial Convolutional Neural Networks (Bi-SCNN) is proposed based on the combination of simplicial convolution with a weighted binary-sign forward propagation strategy. The utilization of the Hodge Laplacian on a weighted binary-sign forward propagation enables Bi-SCNN to efficiently and effectively represent simplicial features with higher-order structures, surpassing the capabilities of traditional graph node representations. The Bi-SCNN achieves reduced model complexity compared to previous SSCN variants through binarization and normalization, also serving as intrinsic nonlinearities of Bi-SCNN; this enables Bi-SCNN to shorten the execution time without compromising prediction performance and makes Bi-SCNN less prone to over-smoothing. Experimenting with real-world citation and ocean-drifter data confirmed that our proposed Bi-SCNN is efficient and accurate.
@@
Graph Learning; Graph Neural Networks; Simplicial Complex; Convolutional Neural Networks; Binarization
@@ 1 Introduction
Graphs and networks are gaining enormous awareness in the field of machine learning and artificial intelligence owing to their ability to effectively learn and represent irregular data interactions (Ortega et al.,, 2018; Leus et al.,, 2023; Sandryhaila and Moura,, 2014). Graph structures can be deduced from various real-life applications by embedding the features onto the nodes of graph representations using the graph Laplacian. Sensor graphs of nationwide temperature (Spelta and Martins,, 2020) and air quality (Wang et al., 2021a, ) are built by combining sensor readings with the sensor location. Citation graphs (McCallum et al.,, 2000) can be formed by linking the citation relationship between authors. The brain region activation can be modeled using a series of dynamic graphs (Zhao and Kuruoglu,, 2024). Graphs and topological structures can be used to model the stock market (Muvunza et al.,, 2024; Qin and Kuruoglu,, 2024). In the past, spatial and spectral graph neural networks (GNNs) such as the Graph Convolutional Networks (GCN) (Defferrard et al.,, 2016; Kipf and Welling,, 2016), Graph Attention Networks (GAT) (Veličković et al.,, 2018), and their variants have had major success on graph representation learning tasks including node classification, graph classification, and link prediction. The GCN and the GAT can be generalized by the Message Passing Neural Networks (MPNN) where instead of Global level graph representation done by the graph Laplacian matrix or the graph adjacency matrix, a localized message-passing scheme is defined for each node (Fey and Lenssen,, 2019). However, GNNs operate only on the features residing on the graph nodes and do not pay enough attention to the graph edges and the structures formed by the edges. In the real world, data or features can appear not only on graph nodes but also on high-order structures. Traffic flow could be recorded on either the edges (Schaub and Segarra,, 2018) or the nodes (Yu et al.,, 2018) of a graph. A map of ocean drifter trajectories can be modeled using the data on graph edges (Schaub et al.,, 2020). In a citation complex, a collaboration among three or even more authors is modeled using simplicial complexes, which is a scenario that cannot be simply recorded using the graph nodes only because a traditional representation of an edge connection between two nodes can only represent a one-to-one relationship between two authors (Ebli et al.,, 2020). More such examples where data appears on structures with high-order than graph nodes include pandemic prediction based on population mobility among regions (Panagopoulos et al.,, 2021) and dynamic gene interaction graphs (Ancherbak et al.,, 2016).
@@
Recently the concept of Topological Signal Processing (TSP) was introduced: instead of the graph Laplacian, a more generalized Laplacian known as the Hodge Laplacian is adopted in TSP to incorporate the representation of high-order features that are beyond the graph nodes (Barbarossa and Sardellitti,, 2020). In TSP, operations such as spectral filtering, spatial diffusion, and convolution are defined on simplicial complexes using the Hodge Laplacians (Yang et al., 2022b, ). High-dimensional feature representations such as data on the graph edges and data on the triangles formed by three edges can be embedded into simplicial complexes using methodologies analogous to graph (node) embedding (Martino et al.,, 2019).
@@
If we feed the outputs of the TSP into non-linear active functions and concatenate a few of these together, we can transform them into neural networks. The Simplicial Neural Network (SNN) uplifts the GCN onto simplicial complexes by using TSP to extend the convolutional ability of GCN to process high-order simplicial features (Ebli et al.,, 2020). The Simplicial Convolutional Neural Networks (SCNN) (Yang et al., 2022a, ) further improves the performance of SNN by applying the Hodge decomposition to the simplicial convolution so it is able to train the upper and lower simplicial adjacencies separately. The Simplicial Attention Network (SAT) (Goh et al.,, 2022) and the Simplicial Attention Neural Network (Giusti et al.,, 2022) introduced the attention mechanism to SNN and SCNN to process simplicial features, making them generalized versions of the GAT from graphs to simplicial complexes. The Simplicial Message Passing Networks (Simplicial MPNN) have made the simplicial complex analogy of defining message passing on graphs (Fey and Lenssen,, 2019) to simplicial complexes (Bodnar et al.,, 2021). The Dist2Cycle was proposed to use Simplicial complexes to define neural networks for Homology Localization (Keros et al.,, 2022). These architectures have shown to be successful at applications such as imputing missing data on the citation complexes of different data dimensions and classifying different types of ocean drifter trajectories.
@@
Although the simplicial complex-based neural networks have achieved numerous successes, the focus of the previous networks was on how to extend concepts such as convolution or attention from graphs onto simplicial complexes, which leads to two drawbacks. First, the recently proposed simplicial neural network structures do not prioritize time efficiency, resulting in inefficiency in terms of computational complexity and run time. Let us look into the GCN, which is the 0-simplex analogy of the SNN, the Simplified GCN (Wu et al.,, 2019) has proved a trained GCN essentially is a low-pass filter in the spectral domain and the complexity of GCN can be reduced by eliminating nonlinearities and collapsing weights. The Binary-GCN (Bi-GCN) (Wang et al., 2021b, ) was inspired by the Binary Neural Networks to use binary operations such as XOR and bit count to speed up training time. In the field of Graph Signal Processing, the Adaptive Graph-Sign algorithm derived from $l_{1}$-optimization results in a simple yet effective adaptive sign-error update that has outstanding run speed and robustness (Yan et al.,, 2022; Yan and Kuruoglu,, 2023). To the best of our knowledge, similar improvements that focus on the efficiency and complexity of neural network algorithms on the simplicial complex are absent. Following the first drawback of the current neural network algorithms on simplicial complexes, the high complexity of current simplicial neural network structures may cause the resulting simplicial embeddings to get over-smoothed in some scenarios. In their graph predecessors, the over-smoothed phenomenon was extensively analyzed (Li et al.,, 2018; Liu et al.,, 2020), with solutions such as spectral filtering (Zhu and Koniusz,, 2021) and regularizations(Chen et al.,, 2023) to alleviate over-smoothing, but has not yet gained enough attention in the simplicial complex-based neural networks.
@@
    The TSP foundation of the Bi-SCNN enables it to effectively represent and process data residing on higher-order simplicial complexes, which extends its learning capabilities beyond just operating on graph nodes.
@@
    The Bi-SCNN can be interpreted using spatial and spectral filters defined using the TSP backbone.
@@
    The combination of simplicial convolution with a weighted binary-sign propagation strategy significantly reduces the model complexity and the training time of Bi-SCNN compared to existing algorithms such as SCNN, without compromising the prediction accuracy. Moreover, the Bi-SCNN inherently incorporates nonlinearity and activation functions through the use of the Sign() function and feature normalization.
@@
    Bi-SCNN is less prone to over-smoothing as a result of the efficient architectural complexity.
This paper is organized as follows. Section 2 provides a discussion of the related work. The preliminary knowledge is discussed in Section 3. The derivation of the Bi-SCNN is presented in Section 4. The experiment results and related discussion can be found in Section 5. Section 6 concludes the paper.
@@ 2 Related Works
Algorithms that conduct convolution on the nodes of a graph such as GCN (Kipf and Welling,, 2016) operate only on the graph nodes, but not on $k$-simplexes with $k\geq 1$. Both the SNN (Ebli et al.,, 2020) and the SCNN (Yang et al., 2022a, ) are direct extensions of the GCN onto simplicial complexes; SNN is architecturally identical to the SCNN except SNN does not use Hodge decomposition. The Bi-SCNN is an extension of the SCNN (Yang et al., 2022a, ), but the Bi-SCNN uses feature normalization and feature binarization which the SCNN does not. Also, the weighted binary-sign forward propagation strategy in the Bi-SCNN layer has 2 outputs but the SCNN layer only has 1 output, making a neural network that contains more than 2 layers of Bi-SCNN a drastically different architecture than using purely SCNN.
These convolutions on graphs or simplicial complexes are enabled using aggregations defined by Laplacian or adjacency matrices where the weights are assigned to the global topology level. To increase the representation power of convolutions on graphs or simplicial complexes, message passing based on lower and upper adjacencies can be defined on the local level (Fey and Lenssen,, 2019; Bodnar et al.,, 2021). On graphs, the MPNN defines trainable weights on all of the edges between the nodes instead of only a small number of weights on the graph (normalized or unnormalized) Laplacian as seen in the GCN (Fey and Lenssen,, 2019). On simplicial complexes, a more powerful but complicated message-passing scheme can be defined using the lower, upper, boundary, and coboundary adjacencies (Bodnar et al.,, 2021). MPNNs indeed boost the representational power of GNNs or SNNs, but this advantage comes with higher memory demands, longer training periods, scalability challenges, and difficulties with over-smoothing due to an increased number of trainable parameters (Di Giovanni et al.,, 2023; Jiang et al.,, 2024).
@@
The Bi-GCN (Wang et al., 2021b, ) has an acronym similar to the Bi-SCNN, but we should emphasize that the Bi-SCNN is not simply a simplicial complex version of the Bi-GCN and they are drastically different from each other with the following four differences. Firstly, the Bi-GCN operates only on the graph nodes, while the Bi-SCNN operates on $k$-simplexes. Secondly, as explained earlier, the Bi-SCNN does not follow the traditional GCN/SCNN forward propagation strategy because the feature binarization and the feature normalization propagate separately. Thirdly, the Bi-GCN binarizes both the feature and the trainable weights but Bi-SCNN only binarizes the features. Fourthly, the Bi-GCN uses binary operations bit count and XOR to speed up the original GCN forward propagation and uses predefined values to deal with the backpropagation of binary operations; the Bi-SCNN uses a weighted binary-sign forward propagation, separating the pass for feature normalization and feature binarization in both forward and backward propagation.
Figure 1: A graph with $N_{0}=24$ nodes, $N_{1}=38$ edges, and $N_{2}=2$ triangles having signals on the edges.
3 Preliminaries
@@
We will begin with a set of vertex $\mathcal{V}={\{v_{0}...v_{N_{0}}}\}$ to define a $k$-simplex $\mathcal{S}_{k}$, where $\mathcal{S}_{k}$ is a subset of $\mathcal{V}$ with cardinality $k+1$. Throughout the paper, subscript $k$ is used to denote which $\mathcal{S}_{k}$ a variable belongs to. The collection of $k$-simplexes with $k=0...K$ is a simplicial complex $\mathcal{K}$ of order $K$. If generalizing a graph $\mathcal{G}=\{\mathcal{V,E}\}$ as a simplicial complex, the vertex set $\mathcal{V}$ of $\mathcal{G}$ is the node set that belongs to $\mathcal{S}_{0}$ and the edge set =$\mathcal{E}=\{e_{1}...e_{N_{1}}\}$ belongs to $\mathcal{S}_{1}$. Note that for a simplex $\mathcal{S}_{k}\in\mathcal{K}$, if $\mathcal{S}_{k-1}$ is a subset of $\mathcal{S}_{k}$, then $\mathcal{S}_{k-1}\in\mathcal{K}$. We use the variable ${N_{k}}$ to denote the number of elements of $\mathcal{S}_{k}$. Take a graph $\mathcal{G}$ as an example, ${N_{0}}=|\mathcal{V}|$ is the number of nodes and ${N_{1}}=|\mathcal{E}|$ is the number of edges. Each simplicial feature vector of a simplex $\mathcal{S}_{k}$ is a mapping from the $k$-simplex to $\mathbb{R}^{N_{k}}$ and will be denoted as $\boldsymbol{x}_{k}=[x_{k,0}...x_{k,N_{k}}]^{T}$ (Barbarossa and Sardellitti,, 2020) . Simplicial data with $d$ features will be denoted as $\mathbf{X}_{k}$ = [$\boldsymbol{x}_{k,0}...\boldsymbol{x}_{k,d}$]. A graph containing only signals on the edges and defined using the above definition is shown in Figure 1.
@@
One can see an important characteristic of $\mathbf{L}_{k}$, for $1\leq k<K$, $\mathbf{L}_{k}$ can be split into $\mathbf{L}_{k,l}$ and $\mathbf{L}_{k,u}$ (1); $\mathbf{L}_{k,l}$ is the lower Hodge Laplacian that represents the lower adjacencies of $\mathcal{S}_{k}$, and $\mathbf{L}_{k,u}$ is the upper Hodge Laplacian that represents the upper-adjacency of $\mathcal{S}_{k}$ (Barbarossa and Sardellitti,, 2020). It is worth pointing out an important property of the simplicial Fourier transform (SFT) in the spectral domain: the eigenvectors paired with the non-zero eigenvalues of $\mathbf{L}_{k,u}$ are orthogonal to the eigenvectors paired with the non-zero eigenvalues of $\mathbf{L}_{k,l}$ (Barbarossa and Sardellitti,, 2020). The graph Laplacian matrix seen in GCN or GAT is $\mathbf{L}_{0}$. Two examples of incidence matrices are $\mathbf{B}_{1}$ and $\mathbf{B}_{2}$, which are the node-to-edge incidence matrix and the edge-to-triangle incidence matrix respectively.
4 Binarized Simplicial Convolutional Neural Networks
@@
We begin with a discussion of spectral and spatial TSP techniques. Then, the simplicial convolution on a $k$-simplex, which is the theoretical foundation of our Bi-SCNN, will be analyzed using TSP. The advantages and drawbacks of several simplicial-convolution-based algorithms are discussed and analyzed. Afterward, we will provide the derivations of Bi-SCNN and combine feature normalization, binary-sign updates, simplicial filters, and nonlinearities to form our Bi-SCNN layer. Lastly, we will give a detailed interpretation of the Bi-SCNN. The goal is to improve the simplicial convolution shown in (9) by reducing its computational complexity and making it less prone to over-smoothing.
4.1 Simplicial Convolution and Topological Signal Processing
In short, the simplicial convolution is essentially the conducting the following procedure: 1. transforming spatial data into the spectral domain, 2. applying a filter in the spectral domain to the simplicial data, 3. using the inverse transform to transform the filtered data back into the spatial domain (Barbarossa and Sardellitti,, 2020). This simple yet efficient procedure is made possible due to the convolution property, which can be summarized as convolution becomes multiplication in the frequency domain. In GNNs, operations such as graph convolution and embedding of the node features $\mathbf{X}_{0}$ are possible by the computations defined by the graph Laplacian. For instance, the spatial graph convolution in GCN aggregates the data from the 1-hop neighbors of each node by $\mathbf{L}_{0}\mathbf{X}_{0}$ (Kipf and Welling,, 2016). From the Graph Signal Processing perspective, the graph convolution can also be defined by applying low-pass filters in the spectral domain to data on the graph nodes (Defferrard et al.,, 2016). In order to define these spectral domain operations on the graph nodes, the node features are viewed as signals and are transformed to the spectral domain using the Graph Fourier Transform. The Graph Fourier Transform is the graph (node) analogy of the Fourier Transform found in classical Signal Processing (Ortega et al.,, 2018). In classical signal processing, signals are projected onto the spectral domain on a Fourier basis. In the case where signals appear on the graph nodes, the eigenvectors of the graph Laplacian matrix $\mathbf{L}_{0}$ serve as the Fourier basis and the corresponding eigenvalues are the frequencies. When the Graph Fourier Transform analogy is made, lower eigenvalues and their corresponding eigenvectors are referred to as the low-frequency components. Intuitively, in the low-frequency components, each signal on the nodes will have similar values to its neighbors; in other words, the low-frequency node signals have smoother variations over the graph topology. On the other hand, higher eigenvalues and their corresponding eigenvectors are the high-frequency components, which represent the signals that change more rapidly over the graph.
In TSP, the SFT is the analogy of the classical Fourier transform, which enabled the definition of spectral operations on simplicial complexes. However, to represent high-order structures, the Hodge Laplacian will be used to define the transform instead of using only the graph Laplacian. For a set of features to be able to transform between domains in a $k$-simplex, we need the SFT defined by the eigendecomposition of the Hodge Laplacian matrix. For the $k$-simplex, the SFT is the eigendecomposition (Barbarossa and Sardellitti,, 2020):